OpenStudy (anonymous):

If you have 0.261 m3 of water at 25.0°C and add 0.127 m3 of water at 95.0°C, what is the final temperature of the mixture? Use 1000 kg/m3 as the density of water at any temperature.

OpenStudy (aaronq):

since they have the same specific heat capacity you can simply write so \(q=m*C_p*\Delta T\) becomes \(q=m*\Delta T\) set them equal to each other and solve for the final temp: \(m_1*(T_f-T_i)=m_2*(T_f-T_i)\) \(0.261\;kg*(T_f-25^oC)=0.127\;kg*(T_f-95^oC)\)